# Improving generalisation via anchor multivariate analysis

### v2

@@ 6.1 Simulation Experiments
6.2 Real Experiment: Robust Climate Prediction
We showcase the efficacy of our approach in a real-world application within the Detection and Attribution of Climate Change (D&A) domain. We extend the methodology of Sippel et al. (2021), who utilised AR for robust detection of forced warming against increased climate variability, by applying it to predict multidimensional local climate responses. Given the increased variability observed in recent climate models Parsons et al. (2020) and observations (DelSole, 2006; Kociuba and Power, 2015; Cheung et al., 2017), our approach aims to ensure robustness against potential underestimation of decadal and multidecadal internal climate variability (Parsons et al., 2020; Mcgregor et al., 2018).
Objective
Specifically, our objective is to use a $p$-dimensional temperature field $X_{\text{mod}}$ to predict a temperature response $Y_{\text{mod}}^{\text{forced}}\in\mathbb{R}^{p}$ to external forcings (such as greenhouse gas emissions, aerosols, solar radiation, or volcanic activity), a crucial step in detecting warming using the fingerprint method (see Hegerl et al., 1996), while ensuring robustness to climate’s Decadal Internal Variability (DIV). To achieve this, we follow Sippel et al. (2021), leveraging multiple climate models from the Climate Model Intercomparison Project (CMIP), both Phase 5 Taylor et al. (2011) and 6 Eyring et al. (2016) (CMIP5 and CMIP6). Specifically, we employ four models (CCSM4, NorCPM1, CESM2, HadCM3) characterised by lower-scale DIV to train our MVA algorithm and validate its robustness using anchor regularisation against models exhibiting higher-scale DIV (CNRM-CM6-1, CNRM-ESM2-1, IPSL-CM6A-LR).
Figure 4: Standard deviation of DIV among training (red) and testing (blue) model members.
Estimators
As a multivariate algorithm, we employ RRRR since both the predictors (temperature fields) and the target (temperature response to external forcing) variables exhibit correlation structures due to spatial autocorrelation. Since both $\alpha$ and $\rho$ in RRRR serve to regularise the regression, we optimise them using cross-validation across the training models. We consider two levels of anchor regularisation: $\gamma=5$ (low regularisation) and $\gamma=100$ to induce high robustness to a potential increase in DIV. We evaluate our results regarding two metrics: $R^{2}$ score (a standard metric when predicting spatial fields) and mean correlation between the anchor (DIV) and the regression residuals (Sippel et al., 2021; Székely et al., 2022).
Data
We selected seven models from the CMIP5 and CMIP6 archives, each containing at least $8$ members from historical simulations, to ensure accurate estimation of the climate response (refer to Tab. 2 for detailed model information). The data preprocessing procedure for each model involves re-gridding surface air temperature data to a regular 5°5  °$\times$5°5  ° grid and computing yearly anomalies by subtracting the mean surface air temperature for the reference period: years $1850-1900$. The forced response $Y^{\text{forced}}_{\text{mod}}$ is obtained using a standard approach (see Deser et al., 2020, and Supplementary details), averaging over all available members in each model. Furthermore, we use DIV as a proxy for multidecadal climate internal variability (see Parsons et al., 2020), achieved by removing the global forced response to global temperature and smoothing it using a $10$-year running mean. This procedure is commonly employed for estimating the multidecadal variability of the climate system (see Sippel et al., 2021; Deser et al., 2020, for further details). Finally, all data are standardised.
Evaluation procedure
We categorise the models into two groups (low and high variability) based on the distribution of the standard deviation of their DIV among the members of each model. As depicted in Fig. 4, a significant proportion of members used for training display a notably smaller magnitude of DIV. Hyperparameters $(\alpha,\rho)$ are selected through a leave-half-models-out cross-validation procedure. Given $k=4$ training models, we randomly select $k/2$ models for training and $k/2$ models for validation, and repeated $B=10$ times. We use $500$ samples from each model to ensure equal weights are given to each training model in the learning algorithm, despite variations in the number of members. For each of the $B$ train/validation splits, we train an RRRR model and select the hyperparameters that yield the highest averaged $R^{2}$ score across the $B$ sampling splits. $R^{2}$ provides a performance measure comparable across regions and is typically used when predicting spatial variables (see e.g. Sippel et al., 2019). Hyperparameter $\alpha$ is selected from $20$ candidates in a logarithmically-spaced sequence ($\lambda\in[1;10^{6}]$), and $\rho$ is selected from $21$ candidates in a linearly-spaced integer sequence ($\rho\in[300;600]$).
@@ Results
For both anchor-regularised RRRR (A-RRRR) with $\gamma=5$ and $\gamma=100$, we observe an improvement in both metrics in the testing set (respectively $0.537$ and $0.533$ of $R^{2}$ for A-RRRR and $0.506$ for unregularised RRRR), while experiencing a slight decrease of performance in the training samples ($0.510$ of $R^{2}$ for unregularised RRRR and respectively $0.500$ and $0.487$ for both A-RRRR). As the latter strongly protects against shifts in DIV, we notice that its mean correlation between residuals and DIV is lower, albeit with a slight decrease in the $R^{2}$ score (see Table 3 in §11). Spatially, we observe (Fig. 5) that anchor-regularised RRRR performs better in the northern hemisphere, particularly in regions where the residuals are highly correlated with DIV and in the northern hemisphere where unregularised RRRR performs poorly (see Figs. 14 and 15 in §11).
Conversely, we observe a decrease in the performance of A-RRRR in regions where unregularised RRRR already performs well or where residuals are anti-correlated with DIV. The latter challenge might be addressed by considering regional proxies of internal variability instead of a global one, which does not always accurately represent local internal variability.
Figure 5: (A) $R^{2}$ score differences between A-RRRR (for $\gamma=5$) and RRRR for testing models. Red values imply that anchor regularisation gives better estimates (larger $R^{2}$ score). (B) Differences of correlation between residuals and DIV ($r$) between anchor-regularised RRRR (for $\gamma=5$) and RRRR for testing models. Red values imply that anchor regularisation gives a lower correlation between residuals and DIV. Regions where A-RRRR perform better are hatched with dots.
7 Conclusion
In this study, we extend the causal framework of anchor regression proposed by Rothenhäusler et al. (2018), demonstrating the versatility of their regularisation approach across a wide range of MVA algorithms. Given the significant challenge of generalising models to OOD data in machine learning, we advocate for integrating anchor regularisation into a broader class of MVA algorithms, including RRR, OPLS, or PLS regression, particularly when prior knowledge is available about the reasons for distribution shifts. Moreover, we highlight that anchor regularisation offers an interesting trade-off between prediction performance and invariance to distribution shifts, addressing concerns regarding over-conservativeness in some cases. Additionally, our application of anchor regularisation in real-world scenarios underscores its practical utility. Future theoretical advancements will entail extending the formulation nonlinear cases using kernel methods Po-Chedley et al. (2022). Furthermore, an intriguing avenue for exploration is understanding how statistical learning algorithms incompatible with anchor regularisation behave when this regularisation is applied. On the application front, we aim to investigate how anchor regularisation, in conjunction with various MVA algorithms, can be leveraged to detect forced responses to more complex climate variables (e.g., precipitation or temperature and their extremes) and attribute them to external forcing sources such as greenhouse gas or aerosol emissions (anthropogenic) or natural phenomena like solar radiation and volcanic activity. We anticipate a broad development and application of the presented and future anchor-regularised methods across various fields of science.
Impact statement
This paper contributes to Machine Learning by showcasing the application of anchor regularisation in both general tasks in multivariate data analysis and statistics and in the specific context of Detection and Attribution (D&A) studies in climate science. The broader impact of this work lies in its potential to enhance the robustness and reliability of algorithms across various domains, including those with societal, economic and environmental implications. Specifically, in climate science, where accurate detection and attribution of climate change impacts are crucial for informed decision-making, our findings demonstrate how anchor regularisation can improve the reliability of predictive models. By addressing distributional shifts and enhancing model robustness, our approach can potentially strengthen the scientific basis for policy-making and adaptation strategies while contributing to the broader understanding and mitigation of climate change impacts.

### v3

@@ D.1 Experiment Setting
We can see that all algorithms exhibit robustness properties for a very large range of perturbation strength. It is interesting to note that CCA presents a similar behavior as the anchor-compatible algorithms. It would be interesting to investigate in which specific cases CCA gives robustness properties.
D.2 Hyperparameters Selection Experiment
Figure 7: Similar experiments as in 6.1 with noise sampled from non-gaussian distributions. (Top) Exponential distribution of scale1111, (Bottom) Poisson distribution of parameter $1$. We see that all the anchor-compatible algorithms are robust to increasing perturbation strength. Shaded areas represent the range of two standard errors of the mean for running $B=20$ times the experiment.
Figure 8: Experiment for anchor confounding $X$ and $Y$. We can see that all algorithms present a very robust behavior to increasing perturbation strength (even CCA which is not anchor-compatible). Shaded areas represent the range of two standard errors of the mean for running $B=20$ times the experiment.
The results of the toy model experiments in the high-dimensional setting are obtained as follows. The training data are sampled following the DAG
$\begin{split}\varepsilon_{A},\,\varepsilon_{H},\,&\varepsilon_{X},\,\varepsilon_{Y}\sim\mathcal{N}(0,1),\\ A&\leftarrow\varepsilon_{A},\\ H&\leftarrow\varepsilon_{H},\\ X&\leftarrow\frac{1}{2}A+H+\varepsilon_{X},\\ Y&\leftarrow 2A+\mathbf{W}^{T}X+H+\varepsilon_{Y},\end{split}$ (26)
and the testing data are generated by modifying $\varepsilon_{A}$ such that $\varepsilon_{A}\sim\mathcal{N}(0,t)$. The perturbation strength $t$ is set to $2$. We repeat each experiment $B=10$ times by resampling $n=100$ training and validation samples and $n=400$ testing samples of $(A,X,Y)$. Both $A$ and $H$ are one-dimensional, while $X$ and $Y$ are $300$-dimensional. The matrix $\mathbf{W}$ is generated as a low-rank (of rank $\rho$ which is randomly sampled uniformly on $[10,30]$) matrix such that $\mathbf{W}=\mathbf{U}\mathbf{V}$ with $\mathbf{U}\in\mathbb{R}^{d\times\rho}$ and $\mathbf{V}\in\mathbb{R}^{\rho\times p}$. The coefficients of $\mathbf{U}$ and $\mathbf{V}$ are sampled uniformly between $1$ and $3$ and are normalised such that their sum is $1$. Hyperparameters $\alpha$ ranging from $1$ to $10^{5}$ are chosen from a logarithmic scale with $20$ candidates, while $\rho$ ranges from $10$ to $30$ linearly with $10$ candidates, and $\gamma$ ranges from $1$ to $10^{4}$ logarithmically with $10$ candidates. Performance, measured in terms of Mean Squared Error and Mean correlation between anchor variable and residuals, is evaluated on a validation set (not seen during training) and a perturbed testing set. We illustrate the results by selecting the optimal rank for each pair $(\gamma,\alpha)$ (see Figure 6). Optimal parameters are chosen to minimize different objectives: for RRRR, parameters $(\alpha,\rho)$ that minimize training MSE are selected, while for A-RRRR, a convex combination of correlation between anchor and residuals ($g_{1}(\gamma,\alpha,\rho)$) and MSE ($g_{2}(\gamma,\alpha,\rho)$) at training is minimized, i.e., $(\gamma,\alpha,\rho)=\operatorname{\text{argmin}}\sum_{i=1}^{2}w_{i}\frac{g_{i}(\gamma,\alpha,\rho)}{\eta_{i}}$, where $\eta_{1}$ and $\eta_{2}$ rescale the two objectives. In Figure 6, we present the results for optimal RRRR and A-RRRR with weights $w_{1}=0.1$ and $w_{2}=0.9$, emphasizing independence of anchor and residuals.
Figure 9: DIV and global forced response for a "low-variability" model (CESM2) (A) and for a "high-variability" model (CNRM-CM6-1) (B).
D.3 Anchor-MVA vs Univariate AR
As mentioned in the section, MVA methods offer a viable solution for handling correlated response variables within the context of the general multivariate linear model $Y=\mu+\mathbf{W}X+\varepsilon$. When $\mathbf{W}$ is not assumed to be full rank, the columns of $Y$ exhibit interdependencies and often demonstrate discernible correlation patterns.
To clarify this point and the extent to which anchor-regularised MVA is advantageous compared to simply running multiple AR, we conducted experiments using a similar setup as in 6.1. We set the intervention strength $t=2$, dimensions $n=200,d=p=400$ with $\mathbf{W}$ of rank $\rho=10$, and used an anchor regularisation parameter $\gamma=5$. We compared single variate AR (run for each output) and anchor-regularised Reduced Rank Regression with an oracle for the rank. The experiment was run $B=50$ times, and we obtained the following results in terms of MSE:
Table 3: Comparaison of multiple AR and A-RRR with $\gamma=5$. Results are reported in term of MSE with 95% confidence intervals.
Multiple AR ($\gamma=5$) A-RRR ($\gamma=5$)
Mean Squared Error $1.91\pm 0.26$ $\mathbf{1.45\pm 0.24}$
This showcases how the use of the anchor-regularised algorithm can outperform AR when faced with high-dimensional output variables.
Appendix E Details of Real-World Experiment
E.1 Robust Climate Prediction
Optimal Fingerprint for Detection of Forced Warming
In D&A studies, the optimal fingerprinting process involves several steps. Firstly, the response of the climate system to an external forcing is extracted using a statistical learning model to predict the forced climate response, denoted as $Y^{\text{forced}}_{\text{mod}}$. This is achieved by utilising spatial predictors from a gridded field of climate variables $X_{\text{mod}}$, following the regression equation:
$Y_{\text{mod}}^{\text{forced}}=X_{\text{mod}}\beta+\beta_{0}+\varepsilon.$ (27)
Here, the spatial fingerprint is represented by the regression coefficient $(\beta,\beta_{0})$. Subsequently, a detection metric is obtained by projecting observations ($X_{\text{obs}}$) and unforced simulations ($X_{\text{cntl}}$) (known as control scenarios) onto the extracted fingerprint $\hat{\beta}$. Practically, as multiple members and models are available for the unforced scenario, a distribution for $X_{\text{mod}}^{\text{cntl}}\hat{\beta}+\hat{\beta}_{0}$ can be obtained. Then, it is tested whether $X_{\text{obs}}\hat{\beta}+\hat{\beta}_{0}$ lies within the same distribution. If the test is rejected, a forced response is detected.
From Global to Regional
Moving from a global to regional scale presents challenges in D&A studies. One of the current challenges is transitioning from detecting a global forced response to a regional or local forced response. This is because climate variability magnitude is larger at regional scales, and even larger at a local scale, leading to a lower signal-to-noise ratio of the forced warming response. As the signal-to-noise ratio is in this case much lower, it becomes much more challenging to detect the forced signal. Additionally, different climate models and observations exhibit different patterns of internal variability at regional scales, necessitating the training of statistical learning models robust to potential distribution shifts in internal variability.
Robustness to Multidecadal Internal Variability
The ability to detect forced warming in Detection and Attribution studies is highly influenced by the level of internal variability (see Parsons et al., 2020). For instance, the observed 40-year Global Mean Temperature (GMT) trend of 0.76°C for the period $1980-2019$ would exceed the standard deviation of natural internal variability in CMIP5 and CMIP6 archive models classified as "low-variability" by a factor of $5$ or more (see Fig. 1B in Sippel et al., 2021). However, for "high-variability" models, this trend exceeds the standard deviation only by a factor of $2$. This becomes evident when comparing models with "low-variability" (Figure 9 A) to those with "high-variability" (Figure 9 B), where it is less clear that internal variability alone could generate the Global Mean Temperature. This discrepancy poses a significant challenge, as in the former case, the observed warming has an extremely low probability of being generated by internal variability alone. Conversely, in the latter case, the rejection of the hypothesis that current temperature trends are solely attributable to internal variability is less robust. For this reason, it is important to develop tests that are robust to changes in the magnitude of internal variability, particularly with the increasing magnitude of DIV. Assuming the DAG depicted in Figure 10, anchor regularisation emerges as a suitable approach, enabling robust estimation robust to DIV increase.
[…]
Figure 10: DAG considered for detection of forced warming response as proposed in Sippel et al. (2021).
Climate Models
In this experiment, we utilised $7$ climate models from the CMIP5 and CMIP6 archives, each having at least $8$ members. The selection of these models was primarily practical, based on our access to them and their suitability for the train/test split procedure based on DIV magnitude.
Table 4: Climate models used in this work, their number of members, associated DIV magnitude and corresponding set (train or test).
[…]
CCSM4 5 8 0.052 Train
NorCMP1 6 30 0.064 Train
CESM2 6 11 0.076 Train
HadCM3 5 10 0.072 Train
CNRM-CM6-1 6 30 0.175 Test
CNRM-ESM2-1 6 10 0.174 Test
IPSL-CM6A-LR 6 32 0.141 Test
@@ Results
Table 5: Metrics for RRRR, and they anchor-regularised version with $\gamma=5$ and $\gamma=100$. We can see that both anchor-regularised approaches outperformed unregularised RRRR in term of $R^{2}$ score and mean correlation of residuals with DIV.
Test Train
$R^{2}$ Mean correlation $R^{2}$ Mean correlation
RRRR 0.506 0.419 0.510 0.157
A-RRRR ($\gamma=5)$ 0.537 0.248 0.500 0.120
A-RRRR ($\gamma=100)$ 0.533 0.123 0.487 0.096
The anchor-regularised version of RRRR generally exhibits superior performance compared to its unregularised counterpart (refer to Table 5 and Figure 12), both in terms of prediction performance (measured via $R^{2}$) and correlation between residuals and DIV at testing time. Notably, these results are achieved at a lower training time cost in terms of $R^{2}$, particularly when utilising $\gamma=5$. This trade-off between performance during training and testing becomes clear when looking at the pareto front emerging (Figure  11) when considering different $\gamma$ values for optimal hyperparameters $(\alpha,\rho)$ selected via the validation procedure detailed earlier. We can see that selecting $\gamma$ such as it minimize an equally weighted ($\gamma 0.5/0.5)$ objectives (prediction performances and correlation between residuals and DIV) lead to a very small decrease during training but induce strong robustness (being close to optimal) at testing in terms of predicting performance.
Regarding the spatial patterns of prediction performance, anchor-regularised RRRR outperforms the unregularised version, particularly in the northern hemisphere and notably at the North Pole, where internal climate variability is significantly higher than closer to the equator. It also exhibits superior performance in northern regions where unregularised RRRR performs poorly, albeit with suboptimal results. Conversely, it demonstrates slightly better performance in the southern hemisphere, especially in regions where the DIV and the residuals exhibit anticorrelation (e.g., central Africa and the southern region of South America). In terms of the correlation between DIV and residuals, anchor-regularised RRRR outperforms the unregularised version across most regions, except for central Africa and the southern regions of South America, where residuals and DIV exhibit anticorrelation. On average, it reduces the correlation between residuals and DIV by more than $15\%$ for $\gamma=5$ and close to $30\%$ for $\gamma=100$.
Figure 11: Pareto front for A-RRRR with optimal $\alpha$, $\rho$, and $\gamma\in[1,10^{2}]$ with optimal $\gamma$ of $/0.5$ 0.5 $0.1/0.9$ weighted objectives.
Figure 12: (Left) $R^{2}$ scores (averaged uniformly). (Right) Mean correlation between DIV and residuals for different values of $\gamma$.
Figure 13: A-RRRR ($\gamma=5$) scores. (A) $R^{2}$ scores. (B) Correlation between DIV and residuals.
Figure 14: A-RRRR ($\gamma=100$) scores. (A) $R^{2}$ scores. (B) Correlation between DIV and residuals.
Figure 15: Unregularised RRRR scores. (A) $R^{2}$ scores. (B) Correlation between DIV and residuals.
E.2 Robust air-quality prediction
IRM and CVP training
IRM and CVP models are trained to MSE convergence with a learning rate of $0.1$, patience of 200 epochs, tolerance of $10^{-4}$, and a maximum of 50,000 epochs.
We provide pseudocodes for both CVP (Alg. 2) and IRM (Alg. 1) algorithms.
Algorithm 1 Invariant Risk Minimization Linear Regression
1:Initialize: Parameters $\phi$, $w$, learning rate $\eta$
2:Input: Training data $X$, $Y$, environment labels $E$, regularization $\lambda_{irm}$, epochs $N$, patience $p$
3:Set best loss to $\infty$, patience counter to 0
4:for $epoch=1$ to $N$ do
5:     Compute IRM loss:
6:      $\mathcal{L}=0,P=0$
7:     for each environment $e$ in $E$ do
8:         Extract $X_{e},Y_{e}$ where $E=e$
9:         Transform input: $X_{e}^{\prime}=X_{e}\phi$
10:         Compute predictions: $\hat{Y}_{e}=X_{e}^{\prime}w$
11:         Compute MSE loss: $\ell_{e}=\text{MSE}(\hat{Y}_{e},Y_{e})$
12:         Compute gradient penalty: $P+=\left(\frac{d\ell_{e}}{dw}\right)^{2}$
13:         Update total loss: $\mathcal{L}+=\ell_{e}$
14:     end for
15:     Compute final loss: $\mathcal{L}=\mathcal{L}+\lambda_{irm}P$
16:     Update $\phi$ and $w$ using gradient descent
17:     Compute validation MSE for early stopping
18:     if validation MSE improves then
19:         Store best $\phi$, $w$
20:         Reset patience counter
21:     else
22:         Increment patience counter
23:     end if
24:     if patience counter $\geq p$ then
25:         Stop training
26:     end if
27:end for
28:Return best $\phi$ and $w$
Algorithm 2 Conditional Variance Penalty Linear Regression
1:Initialize: Model parameters $w$, learning rate $\eta$
2:Input: Training data $X$, $Y$, environment labels $E$, regularization $\lambda_{cvp}$, epochs $N$, patience $p$
3:Set best loss to $\infty$, patience counter to 0
4:for $epoch=1$ to $N$ do
5:     Compute CVP loss:
6:      $\mathcal{L}=\text{MSE}(Xw,Y)$
7:     for each environment $e$ in $E$ do
8:         Extract $X_{e},Y_{e}$ where $E=e$
9:         Compute predictions: $\hat{Y}_{e}=X_{e}w$
10:         Compute variance of predictions: $V_{e}=\text{Var}(\hat{Y}_{e})$
11:         Update total loss: $\mathcal{L}+=\lambda_{cvp}\sum V_{e}$
12:     end for
13:     Update $w$ using gradient descent
14:     Compute validation MSE for early stopping
15:     if validation MSE improves then
16:         Store best $w$
17:         Reset patience counter
18:     else
19:         Increment patience counter
20:     end if
21:     if patience counter $\geq p$ then
22:         Stop training
23:     end if
24:end for
25:Return best $w$
Feature Name Data Type Description
CO(GT) Integer True hourly averaged concentration CO in mg/m3
PT08.S1(CO) Categorical Hourly averaged sensor response
NMHC(GT) Integer True hourly averaged overall Non-Metanic HydroCarbons concentration in µg/m3
C6H6(GT) Continuous True hourly averaged Benzene concentration in µg/m3
PT08.S2(NMHC) Categorical Hourly averaged sensor response
NOx(GT) Integer True hourly averaged NOx concentration in ppb
PT08.S3(NOx) Categorical Hourly averaged sensor response
NO2(GT) Integer True hourly averaged NO2 concentration in µg/m3
Table 6: Air Quality Dataset Outcome Description